One member of the debate team is going to be chosen President. Each member is equally likely to be chosen. The probability that a boy is chosen is 3/4 the probability that a girl is chosen. Girls make up what fraction of the debate team?
Answer by robertb(5830)

Let b = probability that a boy is chosen, and
g = probability that a girl is chosen.
Then from the given,
b = 3g/4.
Also, b + g = 1, because the two probabilities must add up to 1.
By substitution,
3g/4 + g = 1, or 7g/4 = 1, or g = 4/7, the probability that a girl is chosen.
Incidentally, this is also the fraction of girls in the debate team.